Wingate Company, a wholesale distributor of electronic equipment, has been experiencing losses for some time, as shown by its most recent monthly contribution format income statement, which follows:

  

  Sales $ 1,632,000   
  Variable expenses 640,880   
  Contribution margin 991,120   
  Fixed expenses 1,090,000   
  Net operating income (loss) $ (98,880)  

  

In an effort to isolate the problem, the president has asked for an income statement segmented by division. Accordingly, the Accounting Department has developed the following information:

  

Division

East Central West
  Sales $ 422,000 $ 620,000 $ 590,000
  Variable expenses as a percentage of sales 49 % 31 % 41 %
  Traceable fixed expenses $ 251,000 $ 334,000 $ 193,000

Required 1. Prepare a contribution format income statement segmented by divisions, as desired by the president. Division East Central West Total Company
Answer #1
Concepts and reason

Sales revenue: The amount earned by selling goods or services is known as sales revenue. This is the main source of revenue for a company. This is usually referred to as gross sales revenue. The amount of gross sales revenue less returns, allowances, and discounts are known as net sales revenue.

Variable costs: These costs vary with the number of units produced or for the services provided. For example, the labor costs increase if the number of labor hours is increased, and the labor costs decrease if the number of labor hours is decreased.

Fixed costs: These are the costs which remain constant throughout the process of manufacturing or for the services rendered. They are incurred irrespective of number of units produced. For example, factory rent. The rent should be paid, if the factory produces 100 units or 1,000 units or if no units are produced.

Fundamentals

Contribution margin: The difference between the sales revenue and the variable costs is called contribution margin. The following is the formula to compute contribution margin:

Contributionmargin=SalesVariablecosts{\rm{Contribution margin = Sales -- Variable costs}}

Contribution margin ratio: The percentage of sales remained after the variable expenses are paid off is called as contribution margin ratio. The following is the formula to compute contribution margin ratio:

Contributionmarginratio=ContributionmarginSales=SalesVariablecostsSales\begin{array}{c}\\{\rm{Contribution margin ratio}} = \frac{{{\rm{Contribution margin}}}}{{{\rm{Sales}}}}\\\\ = \frac{{{\rm{Sales -- Variable costs}}}}{{{\rm{Sales}}}}\\\end{array}

Income statement: This is the financial statement of a company which reports all the revenues earned and expenses incurred by the company over a specific period of time. Income statement is also known as operations statement, earnings statement, revenue statement, or profit and loss statement.

Contribution format income statements: This income statement is prepared for identifying net operating income in terms of fixed and variable cost behaviors. It helps the managers in decision making, planning and controlling the internal affairs. Here the contribution approach divides cost into two categories as fixed and variable costs.

(1)

Working Notes:

Compute variable expenses of Division East.

Variableexpenses=SalesofDivisionEast×Variableexpensepercentage=$422,000×49%=$206,780\begin{array}{c}\\{\rm{Variable expenses = Sales of Division East \times Variable expense percentage}}\\\\{\rm{ = \$ 422,000 \times 49\% }}\\\\{\rm{ = \$ 206,780}}\\\end{array}

Compute variable expenses of Division Central.

Variableexpenses=SalesofDivisionCentral×Variableexpensepercentage=$620,000×31%=$192,200\begin{array}{c}\\{\rm{Variable expenses = Sales of Division Central \times Variable expense percentage}}\\\\{\rm{ = \$ 620,000 \times 31\% }}\\\\{\rm{ = \$ 192,200}}\\\end{array}

Compute variable expenses of Division West.

Variableexpenses=SalesofDivisionWest×Variableexpensepercentage=$590,000×41%=$241,900\begin{array}{c}\\{\rm{Variable expenses = Sales of Division West \times Variable expense percentage}}\\\\{\rm{ = \$ 590,000 \times 41\% }}\\\\{\rm{ = \$ 241,900}}\\\end{array}

Compute total traceable fixed expenses.

Totaltraceablefixedexpenses}={TraceablefixedexpensesofDivisionEast+TraceablefixedexpensesofDivisionCentral+TraceablefixedexpensesofDivisionWest}=$251,000+$334,000+$193,000=$778,000\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Total traceable }}\\\\{\rm{fixed expenses}}\\\end{array} \right\}{\rm{ = }}\left\{ \begin{array}{l}\\{\rm{Traceable fixed expenses of Division East + }}\\\\{\rm{Traceable fixed expenses of Division Central + }}\\\\{\rm{Traceable fixed expenses of Division West}}\\\end{array} \right\}\\\\{\rm{ = \$ 251,000 + \$ 334,000 + \$ 193,000}}\\\\{\rm{ = \$ 778,000}}\\\end{array}

Compute common fixed expenses not traceable to company.

Commonfixedexpensesnottraceabletodivisions}={TotalfixedexpensesofcompanyTotaltraceablefixedexpenses}=$1,090,000$778,000=$312,000\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Common fixed expenses not }}\\\\{\rm{traceable to divisions}}\\\end{array} \right\}{\rm{ = }}\left\{ \begin{array}{l}\\{\rm{Total fixed expenses of company -- }}\\\\{\rm{Total traceable fixed expenses}}\\\end{array} \right\}\\\\{\rm{ = \$ 1,090,000--\$ 778,000}}\\\\{\rm{ = \$ 312,000}}\\\end{array}

Now, prepare a contribution format income statement of Company W, segmented by divisions as shown below:

Total
Divisions
Particulars
Company
East Central
Sales
$1,632,000 $422.000 $620.000
Less: Variable expenses
640.880 206.780 1

Therefore, divisional segment margin of Division East, Central, and West are $(35,780), $93,800, and $155,100 respectively.

(2-a)

Working Note:

Compute incremental sales of Division West.

Incrementalsales=(SalesofDivisionWest×Incrementalpercentage)=$590,000×19%=$112,100\begin{array}{c}\\{\rm{Incremental sales = }}\left( \begin{array}{l}\\{\rm{Sales of Division West \times }}\\\\{\rm{Incremental percentage}}\\\end{array} \right)\\\\{\rm{ = \$ 590,000 \times 19\% }}\\\\{\rm{ = \$ 112,100}}\\\end{array}

Compute contribution margin ratio.

Contributionmarginratio=ContributionmarginSales=SalesofDivisionWestVariablecostsSalesofDivisionWest=$600,000$246,000$600,000=0.59or59%\begin{array}{c}\\{\rm{Contribution margin ratio}} = \frac{{{\rm{Contribution margin}}}}{{{\rm{Sales}}}}\\\\ = \frac{{{\rm{Sales of Division West -- Variable costs}}}}{{{\rm{Sales of Division West}}}}\\\\ = \frac{{\$ 600,000--\$ 246,000}}{{\$ 600,000}}\\\\{\rm{ = 0}}{\rm{.59 or 59\% }}\\\end{array}

Now, compute incremental net operating income as shown below:

| Computation of Net Operating Income
Incremental sales
$112,100
Contribution margin ratio
x 59%
Incremental contribution mar

Therefore, incremental net operating income is $40,139.

(2-b)

Recommendation: It is recommended to incur additional advertising expenses because the increased advertising expenses would earn additional net operating income (as computed in Part 2-a).

Ans: Part 1

Divisional segment margin of Division East, Central, and West are $(35,780), $93,800, and $155,100 respectively.

Part 2-a

Incremental net operating income is $40,139.

Part 2-b

Yes. It is recommended to increase advertising expenses.